Give two differences that you would observe in the electrolytic cell if copper electrode was made the anode and graphite electrode was made the cathode
1
Expert's answer
2017-11-15T12:56:06-0500
(P1*V1) / T1 = (P2*V2) / T2
756 mmHg = 100791.432 Pa
508 mmHg = 67727.576 Pa
30 C = 303 K
-31 C = 242 K
(100791.432 Pa * 0.33 m3) / 303 K = (V2 * 67727.576 Pa) / 242 K
109.772 = V2*279.866
V2= 2.549 L
Answer:
V= 2.549 L